OpenStudy (anonymous):

The number 0.127 is how much greater than 1/8 ? A. ½ B. 2/10 C. 1/50 D. 1/500 E. 2/500

OpenStudy (anonymous):

First, to compare the two number, they have to be in the same form I'm going to convert 1/8 to a decimal (because its much easier than 0.127 to a fraction) 1/8 in decimal form is .125 .125 _ .127 Since their tenths and hundredths digit are the same you look at their thousandths digit to see which is bigger. 5<7, so .125 < .127 Your answer is D

OpenStudy (anonymous):

\[\frac{127}{1000}-\frac{1}{8}= \frac{1}{500}\]

OpenStudy (anonymous):

the answer should be d 1/500.